When accelaration is varying how to find velocity , direction and displacement.

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Suhas,

In that case we use the following relations,

a= dv/dt = d^2x /dt^2.

and v = dx/dt,

In that case the three equations of motion are invalid !!!!!!

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert